25 pet owners where asked if they live alone or not and what type of pet they had.

Live alone live with others
Cat 3 5
Dog 5 7
Other 2 3

Which 2 events are independent?

A.

other and live alone

B. live alone and live with others

C.

cat and dog

D.

Dog and live alone

E.

Cat and live alone

F. None of the 2 events are independent
Answer #1

P(Others) = (3 + 2)/25 = 1/5

P(Live alone) = (3 + 5 + 2)/25 = 2/5

P(Others and live alone) = 2/25

P(Others)*P(Live alone) = (1/5)*(2/5) = 2/25

Since P(Others)*P(Live alone) = P(Others and live alone), these two events are independent.

Option A is correct.